At constant pressure, a sample of gas occupies 420 mL at 210 K. What volume does the gas occupy at STP?
1
Expert's answer
2020-10-26T14:43:38-0400

V1 / T1 = V2 / T2

V₁ = 420 mL

T₁ = 210 K

V₂ = ?

T₂ = 250 K

420 / 210 = V2 / 250

V2 = (420 / 210) x 250 = 500 mL